OpenStudy (nikki_skellington):

A scale on a map shows that 2 inches = 25 miles. How many inches on the map represent 60 miles?

OpenStudy (jjcoolja12):

nvm

OpenStudy (knov):

if 2 inches = 25miles x inches= 60miles which is 25x=2*60 x=120/25

OpenStudy (nikki_skellington):

ok i got it thx @knov

OpenStudy (knov):

no problem

OpenStudy (nikki_skellington):

could you by any chance help me with one more?

OpenStudy (knov):

k

OpenStudy (nikki_skellington):

How many miles are represented by 1/4 inch on the map?

OpenStudy (nikki_skellington):

it goes along with the first question

OpenStudy (knov):

just like the other one. if 2 inches = 25miles then 1/4 inches= x miles